# Yaohui Li2, Jun Lan1, Huijia Zhu1, Jianfu Zhang3, Weiqiang Wang1,Huaxiong Li2 1Ant Group

### v2

@@ 5.1 Implementation details
$1\times 1$ ACC 100.00 99.86 84.78 95.57 81.43 82.29 100.00 99.79 78.29 81.10 97.64 90.98
AP 93.80 99.64 91.78 96.03 90.04 86.13 99.92 99.81 82.12 86.96 - 92.62
$2\times 2$ ACC 95.71 100.00 98.70 69.14 92.43 93.29 100.00 100.00 83.57 82.94 99.44 92.29
DeMamba AP 85.50 100.00 99.59 76.15 96.78 96.99 99.97 100.00 89.80 89.72 - 93.45
-CLIP-FT $7\times 7$ ACC 98.21 100.00 98.41 65.00 94.00 95.43 100.00 100.00 80.86 84.34 98.19 91.62
AP 87.28 99.99 99.06 73.51 97.47 97.00 99.76 99.97 85.75 89.95 - 92.97
$14\times 14$ ACC 100.00 99.86 93.41 88.57 81.71 70.14 100.00 99.93 76.43 80.02 98.70 89.89
AP 94.94 99.42 97.03 93.56 94.02 84.56 99.12 99.22 77.60 89.46 - 92.89
$1\times 1$ ACC 91.07 100.00 99.35 73.71 89.50 96.29 100.00 100.00 94.71 81.21 99.37 93.20
AP 86.65 99.99 99.83 90.64 97.85 98.81 99.91 99.99 98.43 92.18 - 96.43
$2\times 2$ ACC 98.21 100.00 99.86 65.43 94.86 98.86 100.00 100.00 92.86 89.09 99.42 94.42
DeMamba AP 93.32 100.00 99.97 85.55 98.97 99.60 99.98 100.00 97.77 95.75 - 97.10
-XCLIP-FT $7\times 7$ ACC 96.43 100.00 99.57 56.00 88.71 96.00 100.00 100.00 94.43 79.48 99.64 91.84
AP 96.02 100.00 99.97 82.85 99.69 99.52 99.98 100.00 86.51 97.45 - 96.20
$14\times 14$ ACC 76.79 100.00 99.06 22.86 84.71 95.14 99.52 100.00 87.71 72.25 99.87 85.26
AP 80.43 100.00 99.76 65.69 98.49 99.31 99.79 99.91 96.87 88.31 - 92.86
@@
Model Scan Metric Sora Morph Gen2 HotShot Lavie Show-1 Moon Crafter Model Wild Real Avg.
@@
sweep ACC 95.21 100.00 92.55 61.77 92.69 96.33 99.87 99.32 81.53 73.11 98.79 $90.11_{{\color[rgb]{0,0.453125,0}-3.31}}$
DeMamba AP 93.32 99.95 97.32 85.55 95.88 97.54 99.66 97.54 87.13 80.82 - $93.47_{{\color[rgb]{0,0.453125,0}-3.63}}$
-XCLIP-FT continuous ACC 98.21 100.00 99.86 65.43 94.86 98.86 100.00 100.00 92.86 89.09 99.42 94.42
AP 93.32 100.00 99.97 85.55 98.97 99.60 99.98 100.00 97.77 95.75 - 97.10
@@
Baselines. 1) CLIP (ICML’21) (Radford et al., 2021) is an innovative model that connects images with textual descriptions, enabling it to perform exceptionally well across various vision and language tasks. 2) F3Net (ECCV’20) (Qian et al., 2020) addresses the detection of sophisticated face manipulations using a combination of frequency-aware components and loc al statistics. Its dual-stream approach effectively identifies forgery patterns. 3) VideoMAE (NeurIPS’22) (Tong et al., 2022) adapts the Masked Autoencoder (He et al., 2022) concept for video, enhancing feature extraction and improving performance on downstream tasks through self-supervised training. 4) XCLIP (ECCV’22) (Ni et al., 2022) adapts existing image-language models for video recognition, adding a new cross-frame attention mechanism to improve the exchange of temporal information and tailor prompts specifically for videos. 5) STIL (MM’21) (Gu et al., 2021a) proposes a novel spatial-temporal inconsistency learning method, which captures the spatial-temporal inconsistencies in forged videos by introducing a spatial inconsistency module, a temporal inconsistency module, and an information supplementation module. 6) NPR (CVPR’24) (Tan et al., 2024) is a concept introduced to understand and describe the complex, local interdependencies between image pixels that arise from upsampling processes, particularly evident in synthetic images created by generative models like GANs and diffusion models, which are used to identify and analyze the structural artifacts resulting from these operations.
@@ 5.2 Task1: cross generator generalization
Many-to-many generalization task. This task involves training on 10 baseline categories and then testing on each subset and the average detection performance on $D_{\rm v-ood}$. As shown in Table 3, video models achieve better recognition accuracy compared to image models because video models can model temporal sequences. Moreover, our DeMamba model can be effectively integrated into existing models, achieving significant improvements. For instance, by incorporating the DeMamba module into the XCLIP, the DeMamba-XCLIP-FT attains an average accuracy of 94.42% and an average AP of 97.10%, which represents a 10.22% increase in accuracy and 12.02% increase in AP compared to the original XCLIP. Note that PT (Partially Tuning) indicates that the backbone is frozen, with only the other parts being fine-tuned, while FT (Full Training) tuning the entire model.
@@ Appendix A Model details
Model Detection level Input size Param (M) FLOPs(G) ACC AP
CLIP-B Image $8\times 224\times 224$ 151.46 149.34 91.19 91.52
NPR Image $8\times 224\times 224$ 1.44 14.08 85.30 82.45
F3Net Image $8\times 299\times 299$ 48.31 145.04 83.45 88.73
VideoMAE-B Video $16\times 224\times 224$ 86.54 147.52 88.61 94.54
STIL Video $8\times 224\times 224$ 22.69 38.06 83.71 87.12
XCLIP-B Video $8\times 224\times 224$ 121.26 141.28 85.66 86.77
DeMamba-XCLIP Video $8\times 224\times 224$ 125.37 147.61 94.42 97.10

### v3

@@ 5.1 Implementation details
$1\times 1$ R 1.0000 0.9986 0.8478 0.9557 0.8143 0.8229 1.0000 0.9979 0.7829 0.8110 0.9031
F1 0.3218 0.8556 0.8398 0.8334 0.8216 0.7624 0.8414 0.9215 0.7390 0.7849 0.7722
AP 0.9380 0.9964 0.9178 0.9603 0.9004 0.8613 0.9992 0.9981 0.8212 0.8696 0.9262
$2\times 2$ R 0.9571 1.0000 0.9870 0.6914 0.9243 0.9329 1.0000 1.0000 0.8357 0.8294 0.9158
F1 0.6463 0.9615 0.9739 0.7803 0.9414 0.9276 0.9572 0.9804 0.8723 0.8782 0.8919
DeMamba AP 0.8550 1.0000 0.9959 0.7615 0.9678 0.9699 0.9997 1.0000 0.8980 0.8972 0.9345
-CLIP-FT $7\times 7$ R 0.9821 1.0000 0.9841 0.6500 0.9400 0.9543 1.0000 1.0000 0.8086 0.8434 0.9163
F1 0.3724 0.8855 0.9308 0.6811 0.9085 0.8630 0.8737 0.9393 0.7828 0.8271 0.8064
AP 0.8728 0.9999 0.9906 0.7351 0.9747 0.9700 0.9976 0.9997 0.8575 0.8995 0.9297
$14\times 14$ R 1.0000 0.9986 0.9341 0.8857 0.8171 0.7014 1.0000 0.9993 0.7643 0.8002 0.8901
F1 0.4628 0.9149 0.9214 0.8550 0.8555 0.7430 0.9059 0.9556 0.7845 0.8245 0.8223
AP 0.9494 0.9942 0.9703 0.9356 0.9402 0.8456 0.9912 0.9922 0.7760 0.8946 0.9289
$1\times 1$ R 0.9107 1.0000 0.9935 0.7371 0.8950 0.9629 1.0000 1.0000 0.9471 0.8121 0.9258
F1 0.5952 0.9569 0.9748 0.8066 0.9227 0.9387 0.9521 0.9780 0.9298 0.8644 0.8919
AP 0.8665 0.9999 0.9983 0.9064 0.9785 0.9881 0.9991 0.9999 0.9843 0.9218 0.9643
$2\times 2$ R 0.9821 1.0000 0.9986 0.6543 0.9486 0.9886 1.0000 1.0000 0.9286 0.8909 0.9392
F1 0.6467 0.9602 0.9790 0.7539 0.9537 0.9551 0.9557 0.9797 0.9240 0.9120 0.9020
DeMamba AP 0.9332 1.0000 0.9997 0.8555 0.9897 0.9960 0.9998 1.0000 0.9777 0.9575 0.9710
-XCLIP-FT $7\times 7$ R 0.9643 1.0000 0.9957 0.5600 0.8871 0.9600 1.0000 1.0000 0.9443 0.7948 0.9106
F1 0.7013 0.9689 0.9821 0.6895 0.9244 0.9485 0.9653 0.9842 0.9409 0.8622 0.8967
AP 0.9602 1.0000 0.9997 0.8285 0.9969 0.9952 0.9998 1.0000 0.8651 0.9745 0.9620
$14\times 14$ R 0.7679 1.0000 0.9906 0.2286 0.8471 0.9514 0.9952 1.0000 0.8771 0.7225 0.8380
F1 0.7748 0.9908 0.9909 0.3670 0.9126 0.9659 0.9873 0.9954 0.9253 0.8326 0.8742
AP 0.8043 1.0000 0.9976 0.6569 0.9849 0.9931 0.9979 0.9991 0.9687 0.8831 0.9286
@@
Model Scan Metric Sora Morph Gen2 HotShot Lavie Show-1 Moon Crafter Model Wild Avg.
@@
sweep R 0.9521 1.0000 0.9255 0.6177 0.9269 0.9633 0.9987 0.9932 0.8153 0.7311 0.8924
F1 0.4649 0.9211 0.9200 0.6906 0.9212 0.9029 0.9124 0.9557 0.8207 0.7856 $0.8295_{{\color[rgb]{0,0.453125,0}-0.0725}}$
DeMamba AP 0.9332 0.9995 0.9732 0.8555 0.9588 0.9754 0.9966 0.9754 0.8713 0.8082 $0.9347_{{\color[rgb]{0,0.453125,0}-0.0363}}$
-XCLIP-FT continuous R 0.9821 1.0000 0.9986 0.6543 0.9486 0.9886 1.0000 1.0000 0.9286 0.8909 0.9392
F1 0.6467 0.9602 0.9790 0.7539 0.9537 0.9551 0.9557 0.9797 0.9240 0.9120 0.9020
AP 0.9332 1.0000 0.9997 0.8555 0.9897 0.9960 0.9998 1.0000 0.9777 0.9575 0.9710
@@
Baselines. Three type of image-level methods and six video-level methods are treated as baselines for comparison: F3Net Qian et al. (2020), NPR (Tan et al., 2024), CLIP (Radford et al., 2021), STIL (Gu et al., 2021a), XCLIP (Ni et al., 2022), VideoMAE (Tong et al., 2022), MINTIME (Coccomini et al., 2024), FTCN (Zheng et al., 2021), TALL (Xu et al., 2023a).
@@ 5.2 Task1: cross generator generalization
Many-to-many generalization task. This task involves training on 10 baseline categories and then testing on each subset and the average detection performance on $D_{\rm v-ood}$. As shown in Table 3, video models achieve better recognition accuracy compared to image models because video models can model temporal sequences. Moreover, our DeMamba model can be effectively integrated into existing models, achieving significant improvements. For example, integrating the DeMamba module into XCLIP results in DeMamba-XCLIP-FT achieving an average recall/F1/AP of 0.9392/0.9020/0.9710 respectively, which marks an improvement of 11.26%/17.72%/12.02% in recall/F1/AP over the original XCLIP. Note that PT (Partially Tuning) indicates that the backbone is frozen, with only the other parts being fine-tuned, while FT (Full Training) tuning the entire model.
@@ 5.4 Ablation study
Figure 3: Performance of training on scaled-up datasets on the testing set.
Necessity of large volume of GenVideo for generalizability. To clarify the impact of data volume on detection performance, we conducted experiments using DeMamba on various scaled subsets of the GenVideo dataset. We designed two experiments to explore: 1) variation in the number of training samples within each generator: For the data generated by each generator, we randomly selected between 100 and 20,000 video samples to train the model. 2) variation in the number of training samples across different generators: We randomly selected subsets to simulate scenarios with 1 to 10 generators in the dataset. In each scenario, we collected 20,000 video samples for the videos generated by each generator. Note that, for real videos, we randomly selected an equal number of samples to the total number of training generated video data to participate in the training. The test sets used in these experiments are the same as those mentioned in the main paper, they consist of the full test set videos, and the model selected for the experiment is DeMamba-XCLIP-FT. The comparative results of these experiments are detailed in Figure 4. The results clearly indicate a significant association between the scale of the dataset and the improvement in detection performance, further affirming the critical role of massive data in enhancing detection capabilities.
Table 9: Training/inference time and QPS for different models.
Model Training time (h) Inference time (s) QPS
F3Net 40 188 2.47
NPR 25.1 149 2.66
STIL 11.7 174 2.42
VideoMAE 52.5 283 1.32
MINTIME 18.1 148 2.61
FTCN 11.8 148 2.68
TALL 6 150 2.71
CLIP-B-FT 10.1 148 2.56
DeMamba-CLIP-B-FT 14 150 2.54
XCLIP-FT 11.5 149 2.55
DeMamba-XCLIP-FT 15.1 153 2.53
Training and inference efficiency. As detailed in Table 9, we report the training and inference times on 8 Tesla A100-80G GPUs, with QPS measured on a single GPU at batch size 1. Incorporating DeMamba into XCLIP does not significantly impact inference time compared to CLIP and XCLIP. Slower training is due to Mamba’s training characteristics, but it offers efficient inference post-training. Our method maintains competitive computational times relative to other video models, outperforming F3Net, STIL, and TALL.
@@ Appendix A Model details
Model Detection level Input size Param (M) FLOPs(G) R F1 AP
CLIP-B Image $8\times 224\times 224$ 151.46 149.34 0.9057 0.7631 0.9152
NPR Image $8\times 224\times 224$ 1.44 14.08 0.8408 0.7164 0.8245
F3Net Image $8\times 299\times 299$ 48.31 145.04 0.8188 0.8024 0.8873
VideoMAE-B Video $16\times 224\times 224$ 86.54 147.52 0.8750 0.8955 0.9454
STIL Video $8\times 224\times 224$ 22.69 38.06 0.8222 0.7823 0.8712
MINTIME-CLIP-B Video $8\times 224\times 224$ 201.35 221.0 0.8762 0.8231 0.9155
FTCN-CLIP-B Video $8\times 224\times 224$ 174.82 146.69 0.8618 0.8755 0.9221
TALL Video $8\times 224\times 224$ 86.92 15.2 0.8852 0.7419 0.8791
XCLIP-B Video $8\times 224\times 224$ 121.26 141.28 0.6615 0.6497 0.7154
DeMamba-XCLIP Video $8\times 224\times 224$ 125.37 147.61 0.9392 0.9020 0.9710